OpenStudy (anonymous):

A 0.100-kilogram apple falls from a height of 1.50 meters to 1.00 meter. Ignoring frictional effects, the total mechanical energy of the apple at this height is _____. 0.00 J less than 1.47 J but more than 0.00 J exactly 1.47 J more than 1.47 J

OpenStudy (anonymous):

Mechanical energy consists of potential energy and kinetic energy. Potential Energy: \[PE = m \times g \times h\] Kinetic Energy: \[KE = \frac{ 1 }{ 2 }mv^2\] Where, m = mass g = gravity h = height v = velocity

OpenStudy (anonymous):

So, what's the answer?

OpenStudy (anonymous):

Use conservation of energy principles to solve the problem

OpenStudy (anonymous):

more than 1.47 J?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Smh nevermind